Perform the indicated operations and simplify your answer.
a) (5x - y) + (4x + 7)
b) (x^3 + x^2 - 10x - 6) - (x - 3)
c) (-4x^3)^2
d) 8x(x - 2)
e) (5x - 3)(4x - 7)
f) (p^3 + 3)^2
g) (x^2 + 5)(x^2 + 4)
h) (6y - 1)(5y + 1)

Answer by papugai(18)

a) (5x - y) + (4x + 7) = 9x-y+7
b) (x^3 + x^2 - 10x - 6) - (x - 3) = x^3 + x^2 - 11x - 3
c) (-4x^3)^2 = 16x^6
d) 8x(x - 2)= 8x^2 - 16x
e) (5x - 3)(4x - 7) = 20x -47x + 21
f) (p^3 + 3)^2 = p^6 +6p^3 + 9
g) (x^2 + 5)(x^2 + 4) = x^4 + 9x^2 +20
h) (6y - 1)(5y + 1) = 30y^2 ++y -1
These are the answers. Please tell me if you need me to go step by step through the questions.